# Improving Out-of-Distribution Generalization in Graphs via Hierarchical Semantic Environments

### v1

@@ 4.1 Hierarchical Stochastic Subgraph Generation
Given a graph $G=(V,E)$, where $V$ represents the set of nodes and $E$ represents the set of edges, the goal is to generate the invariant subgraph $G^{k}_{v}$ and variant subgraph $G^{k}_{inv}$ from the original graph $G$ at each hierarchy $k\in K$, where $K$ denotes the number of environmental hierarchies. At each hierarchy $k$, we employ a graph neural network denoted as $\text{GNN}_{k}$ to update the hidden embedding $h^{k}_{i}=\text{GNN}_{k}(h^{k-1}_{i},A)$ for each node $v_{i}\in V$. We use graph neural networks as $\text{GNN}_{k}$ [42] to aggregate information from neighboring nodes for node embedding updating. In the process of learning variant and invariant subgraphs at each hierarchy, we define a probability distribution function $S^{k}(\cdot)$ for edge selection as follows. We omit the hierarchy notations, while the computation of probability distribution is conducted within each hierarchy:
@@ 6 Conclusion
In our research, we take on the formidable challenge of improving out-of-distribution (OOD) generalization in the graph domain. We observe that recent works often overlook the crucial aspect of studying hierarchical environments in graph invariant learning. To address this gap, we introduce a novel method that generates hierarchical dynamic environments for each graph. Our approach involves hierarchical stochastic subgraph generation, hierarchical environment inference, and a carefully designed learning objective. By incorporating graph invariant learning with inferred high-level environments, our model not only achieves meaningful and diverse environments within the same hierarchy but also ensures consistency across different hierarchies. The effectiveness of our method is particularly pronounced in the DrugOOD dataset, shedding light on the potential for further exploration in hierarchical graph learning within OOD scenarios.

### v2

@@ 4.1 Hierarchical Stochastic Subgraph Generation
Given a graph $G=(V,E)$, where $V$ represents the set of nodes and $E$ represents the set of edges, the goal is to generate the variant subgraph $G^{k}_{v}$ and invariant subgraph $G^{k}_{inv}$ from the original graph $G$ at each hierarchy $k\in K$, where $K$ denotes the number of environmental hierarchies. At each hierarchy $k$, we employ a graph neural network denoted as $\text{GNN}_{k}$ to update the hidden embedding $h^{k}_{i}=\text{GNN}_{k}(h^{k-1}_{i},A)$ for each node $v_{i}\in V$. We use graph neural networks as $\text{GNN}_{k}$ [42] to aggregate information from neighboring nodes for node embedding updating. In the process of learning variant and invariant subgraphs at each hierarchy, we define a probability distribution function $S^{k}(\cdot)$ for edge selection as follows. We omit the hierarchy notations, while the computation of probability distribution is conducted within each hierarchy:
@@ 6 Conclusion
In our research, we take on the formidable challenge of improving out-of-distribution (OOD) generalization in the graph domain. We observe that recent works often overlook the crucial aspect of studying hierarchical environments in graph invariant learning. To address this gap, we introduce a novel method that generates hierarchical dynamic environments for each graph. Our approach involves hierarchical stochastic subgraph generation, hierarchical environment inference, and a carefully designed learning objective. By incorporating graph invariant learning with inferred high-level environments, our model achieves meaningful and diverse environments within the same hierarchy and ensures consistency across different hierarchies. The effectiveness of our method is particularly pronounced in the DrugOOD dataset, shedding light on the potential for further exploration in hierarchical graph learning within OOD scenarios.
7 Acknowledgement
This work was supported by Bio & Medical Technology Development Program of the NRF & ICT(NRF-2022M3E5F3085677, RS-2023-00257479), RS-2023-00246586 and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [RS-2021-II211343]